OpenStudy (anonymous):

The graph of which quadratic equation is shown below? y = x2 + 4x − 3 y = x2 − 4x + 3 y = −x2 + 4x − 3 y = −x2 − 4x + 3

OpenStudy (anonymous):

where is the graph??

OpenStudy (anonymous):

sorry i thought it posted

OpenStudy (anonymous):

option B is correct

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

general form of quadratic equation is ax2+bx+c.if a>0,graph is an upward parabola,c determines the point where the graph meets the y axis.in this case,we have an upward parabola,so a=3,and the graph cuts y axis at y=3,therefore c=3

OpenStudy (jack1):

easiest way to check your answer is to sub in some figures for x (both positive and negative) and see what corresponding y values it gives you, then plot to see which matches

OpenStudy (anonymous):

thanks. can u help with another ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Tom wants to hang a mirror in his room but the mirror and frame must not have an area larger than 14 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.) 4x2 + 14x − 2 = 0 3x2 + 10x − 14 = 0 2x2 + 12x − 14 = 0 x2 + 7x − 14 = 0

OpenStudy (anonymous):

please check the first option,i think it should be 12 instead of 2

OpenStudy (anonymous):

well it was given to me as 2 so idk

OpenStudy (anonymous):

 basicallicy theb question is asking us to calculate the area of the rectangle ABCD,for that we need its length AD and breadth AB.the area is then AB*AD